OpenStudy (anonymous):

All real numbers greater than or equal to 38.

OpenStudy (anonymous):

What exactly are you trying to find out?

OpenStudy (anonymous):

How to write this as an equality. Sorry I didn't specify that.

OpenStudy (anonymous):

The number of points scored in a basketball game by Slim Dunkin is shown in the table at right: a) What is the probability that he scores less than 11 points in a game? b) What is the probability that he scores over 20 points in a game? c) Which type of probability is this? Give a reason for your answer. d) Is this sample space discrete or continuous? Explain. points per game frequency 6-10 1 11-15 6 16-20 5 21- 25 17 25+ 8

OpenStudy (anonymous):

Ok. If x is your variable...\[x \ge 38\]

OpenStudy (anonymous):

i didn't get it

OpenStudy (anonymous):

That was a reply to Jeremiah's question.

OpenStudy (anonymous):

ok can u help me for my question

OpenStudy (anonymous):

I would need to see the mentioned table. But you should post it in your own question thread.

OpenStudy (anonymous):

The number of points scored in a basketball game by Slim Dunkin is shown in the table at right: a) What is the probability that he scores less than 11 points in a game? b) What is the probability that he scores over 20 points in a game? c) Which type of probability is this? Give a reason for your answer. d) Is this sample space discrete or continuous explain

OpenStudy (anonymous):

points per game frequency 6-10 1 11-15 6 16-20 5 21- 25 17 25+ 8

OpenStudy (anonymous):

a. Probability of less than 11 / Total =\[\frac{ 1 }{ 1+6+5+17+8 }\] b. Probability of 20+ / Total = \[\frac{ 17+8 }{ 1+6+5+17+8 }\]

OpenStudy (anonymous):

thnx so much

OpenStudy (anonymous):

what about question c & d

OpenStudy (anonymous):

I'm not sure. If you still need help after I finish dinner I'll see if I can look it up for you.

OpenStudy (anonymous):

ok that's will be great

OpenStudy (anonymous):

did u have ur dinner i'm waiting

OpenStudy (anonymous):

when u done can u respond my question